# Who Should I Listen To? Adaptive Collaboration in Personalized Federated Learning

### v2

@@ A Note on the Byzantine Resilience of FedMosaic
Following the argument by [Jiang et al., 2020], who show that federated semi-supervised learning with soft labels sharing (e.g., FedDistill) is more Byzantine resilient than FedAvg due to the bounded nature of the threat vector on the probability simplex, we argue that FedMosaic exhibits similar ( if not stronger) resilience properties. Like FedCT [Abourayya et al., 2025], FedMosaic relies on hard label sharing, further constraining the threat vector to a binary classification decision per example. Moreover, FedMosaic incorporates confidence-based aggregation, which naturally downweights unreliable predictions. This mechanism provides an additional layer of robustness by reducing the influence of low confidence ( and potentially malicious) clients. While a formal analysis remains open, these properties suggest that FedMosaic may be at least as Byzantine resilient as FedDistill and FedCT. Exploring this direction further is promising for future work.
8 Details on Experiments
@@
FashionMinst and CIFAR10 datasets have been used for label skew experiments. In FashionMinst, we converted the raw grayscale $28\times 28$ images into Pytorch tensors and normalized pixel values to the range $[-1,1]$ using a mean of $0.5$ and standard deviation of $0.5$. In CIFAR10, we converted RGB $32\times 32$ images into Pytorch tensors of shape $[3,32,32]$ and normalizes each color channel independently to the range of $[-1,1]$, using a mean of $0.5$ and standard deviation of $0.5$. The data is partitioned across $15$ clients. In a pathological non-IID setting, each client receives data from only $2$ out of $10$ classes. In a practical non-IID setting, data is distributed across $15$ clients using a Dirichlet distribution. This creates naturally overlapping, imbalanced label distributions among clients. Training data distribution of each scenario of CIFAR10 are showing in Fig.7 and Fig.7.
Figure 6: CIFAR10 clients data distribution
@@
Figure 7: CIFAR10 clients data distribution
@@
we used the Office-Caltech10 and DominNet datasets. For both, we adopt AlexNet as a neural network architecture. Input images are resized to $256\times 256\times 3$. Training is performed till convergence using the corss-entropy loss and Adam optimizer with learning rate of $10^{-2}$. We use a batch size of $32$ for Office-Caltech10 dataset and $64$ for DomainNet. For DomainNet, which originally contains $345$ categories, we restrict the label space to the top $10$ most frequent classes to reduce complexity, The selected categories are: bird, feather, headphones, icecream, teapot, tiger, whale, windmill, wineglass, zebra. For Office-Caltech10, each client get one of the $4$ domains and For DomainNet dataset, each client get one of the $6$ domains. The distribution of each client training data are showing in Fig.9 and Fig.9.
@@
Figure 9: Office-Caltech10 clients data distribution.
@@
We simulate the hybrid data distribution by combining both label distribution skew and feature distribution shift. We use the same two datasets as in feature shift experiments: Office-Caltech10 and DomainNet. To introduce label skew, for each domain, we randomly sample $5$ clients and assign to each client only $2$ out of $10$ total classes. This results in $20$ clients for the Office-Caltech10 dataset ($4$ domains $\times$ $5$ clients) and $30$ clients for DomainNet ($6$ domains $\times$ $5$ clients). This creates a hybrid non-IID setting where clients differ significantly in both input distribution and output distribution. We use the same preprocessing and training configurations as the feature shift experiments. All input images are resized to $256\times 256\times 3$ before being fed into $AlexNet$. Models are trained using cross-entropy loss and Adam optimizer with learning rate of $10^{-2}$. The batch size is set to $32$ for Office-Caltech10 and $64$ for DomainNet. For DomainNet, we selected the $10$ most frequent as feature shift experiments.To effectively visualize the distribution of local training data across $30$ clients, we used a dot matrix plot, which offers a compact and intuitive representation of client-level variation. The visualization of the Clients distribution of DomainNet and Office-Caltech10 datasets are shown in Fig.11 and Fig.11
@@
Figure 11: Office-Caltech10 clients Hybrid data distribution.
9 Practical Impact of FedMosaic
FedMosaic demonstrates substantial practical impacts in personalized federated learning (PFL), specifically addressing data heterogeneity across decentralized clients. FedMosaic uses a fine-grained collaboration mechanism that enables each client to selectively rely on the collective expertise available, enhancing model accuracy and robustness. This approach is particularly valuable in domains with significant data variability, such as healthcare, finance, and personalized recommendation systems, where traditional federated methods often fail. Empirical results highlight that FedMosaic consistently outperforms state-of-the-art PFL methods, as well as local and centralized baselines across various challenging scenarios, including label skew, feature shift, and hybrid heterogeneity distribution. Its design inherently preserves privacy by limiting data exchanges to hard predictions over a shared unlabeled dataset, thus minimizing potential privacy leakage compared to parameter sharing approaches. This follows the approach to share as little as possible [Mian et al., 2023, Tan et al., 2022] and therefore adheres to the privacy-by-design-principle [Cavoukian et al., 2009]. Furthermore, FedMosaic’s differentially private variant (DP-FedMosaic) showcases the feasibility of achieving robust privacy guarantees without significant compromise on performance, making it particularly attractive for sensitive applications. Another aspect is the communication efficiency of federated co-training approaches for large models: in those cases the number of parameters vastly exceed the number of labels for the public dataset. Combining this with communication-efficient protocols [Kamp et al., 2016, Kamp, 2019] has the potential to reduce communication by several orders of magnitude, in particular for large transformer-based models, such as LLMs.

### v3

@@ A Note on the Byzantine Resilience of FedMosaic
Following the argument by (Jiang et al., 2020), who show that federated semi-supervised learning with soft labels sharing (e.g., FedDistill) is more Byzantine resilient than FedAvg due to the bounded nature of the threat vector on the probability simplex, we argue that FedMosaic exhibits similar ( if not stronger) resilience properties. Like FedCT (Abourayya et al., 2025), FedMosaic relies on hard label sharing, further constraining the threat vector to a binary classification decision per example. Moreover, FedMosaic incorporates confidence-based aggregation, which naturally downweights unreliable predictions. This mechanism provides an additional layer of robustness by reducing the influence of low confidence ( and potentially malicious) clients. While a formal analysis remains open, these properties suggest that FedMosaic may be at least as Byzantine resilient as FedDistill and FedCT. Exploring this direction further is promising for future work.
Appendix C Details on Experiments
All experiments are conducted for a sufficient number of communication rounds until convergence, using three different random seeds. While the standard deviation across the three runs with different seeds is consistently small, this observation aligns with prior work Zhang et al. (2023d),Zhang et al. (2023c), Zhang et al. (2023b).
@@ Appendix C Details on Experiments
Fashion-Minst and CIFAR-10 datasets have been used for label skew experiments. In Fashion-Minst, we converted the raw grayscale $28\times 28$ images into Pytorch tensors and normalized pixel values to the range [1,1][-1,1] using a mean of $0.5$ and standard deviation of $0.5$. In CIFAR-10, we converted RGB $32\times 32$ images into Pytorch tensors of shape $[3,32,32]$ and normalizes each color channel independently to the range of [1,1][-1,1], using a mean of $0.5$ and standard deviation of $0.5$. The data is partitioned across $15$ clients. In a pathological non-IID setting, each client receives data from only $2$ out of $10$ classes. In a practical non-IID setting, data is distributed across $15$ clients using a Dirichlet distribution. This creates naturally overlapping, imbalanced label distributions among clients. Training data distribution of each scenario of CIFAR-10 are showing in Fig.12 and Fig.12.
Figure 11: CIFAR-10 clients data distribution
@@
Figure 12: CIFAR-10 clients data distribution
@@
we used the Office-10 and DomainNet datasets. For both, we adopt AlexNet as a neural network architecture. Input images are resized to $256\times 256\times 3$. Training is performed till convergence using the corss-entropy loss and Adam optimizer with learning rate of $10^{-2}$. We use a batch size of $32$ for Office-10 dataset and $64$ for DomainNet. For DomainNet, which originally contains $345$ categories, we restrict the label space to the top $10$ most frequent classes to reduce complexity, The selected categories are: bird, feather, headphones, icecream, teapot, tiger, whale, windmill, wineglass, zebra. For Office-10, each client get one of the $4$ domains and For DomainNet dataset, each client get one of the $6$ domains. The distribution of each client training data are showing in Fig.14 and Fig.14.
@@
Figure 14: Office-10 clients data distribution.
@@
We simulate the hybrid data distribution by combining both label distribution skew and feature distribution shift. We use the same two datasets as in feature shift experiments: Office-10 and DomainNet. To introduce label skew, for each domain, we randomly sample $5$ clients and assign to each client only $2$ out of $10$ total classes. This results in $20$ clients for the Office-Caltech10 dataset ($4$ domains $\times$ 55 clients) and $30$ clients for DomainNet ($6$ domains $\times$ 55 clients). This creates a hybrid non-IID setting where clients differ significantly in both input distribution and output distribution. We use the same preprocessing and training configurations as the feature shift experiments. All input images are resized to $256\times 256\times 3$ before being fed into $AlexNet$. Models are trained using cross-entropy loss and Adam optimizer with learning rate of $10^{-2}$. The batch size is set to $32$ for Office-10 and $64$ for DomainNet. For DomainNet, we selected the $10$ most frequent as feature shift experiments.To effectively visualize the distribution of local training data across $30$ clients, we used a dot matrix plot, which offers a compact and intuitive representation of client-level variation. The visualization of the Clients distribution of DomainNet and Office-10 datasets are shown in Fig.16 and Fig.16
@@
Figure 16: Office-10 clients Hybrid data distribution.
Appendix D Practical Impact of FedMosaic
FedMosaic addresses data heterogeneity in personalized federated learning (PFL) via a fine-grained collaboration mechanism that lets each client selectively rely on collective expertise, aiming to improve accuracy and robustness. This is particularly relevant in domains with substantial variability (e.g., healthcare, finance, recommendation), where traditional federated methods can struggle. Empirically, FedMosaic often outperforms strong PFL baselines and, in our evaluated settings, local and centralized training across label skew, feature shift, and hybrid heterogeneity; where margins are small, it performs comparably. Its design limits disclosure by sharing only hard predictions on a shared unlabeled dataset, reducing potential privacy leakage relative to parameter sharing. This follows “share as little as possible” (Mian et al., 2023; Tan et al., 2022) and aligns with privacy-by-design (Cavoukian et al., 2009). In addition, our differentially private variant (DP-FedMosaic) illustrates how to obtain formal $(\varepsilon,\delta)$-DP guarantees for the released signals (labels and expertise), with the privacy accounting provided and empirical calibration left to future work. Finally, federated co-training is communication-efficient for large models: when parameter counts vastly exceed $|U|$, sending hard labels (and one expertise scalar per example) can reduce uplink by orders of magnitude. Combining this with communication-efficient protocols (Kamp et al., 2016; Kamp, 2019) has the potential to reduce communication by several orders of magnitude, in particular for large transformer-based models, such as LLMs.
Appendix E Notation
Federated Learning Setup
$m$ Number of participating clients
$i\in[m]$ Index of a client
$D_{i}$ Private dataset of client $i$
$U$ Shared public unlabeled dataset used for co-training
$T$ Total number of communication rounds
$b$ Communication period (local steps between rounds)
$A_{i}$ Local learning algorithm used by client $i$
Models and Predictions
$h_{i}^{t}$ Local model of client $i$ at round $t$
$L(h,D)$ Loss of model $h$ on dataset $D$
$\ell_{\text{priv}}=L(h_{i}^{t-1},D_{i})$ Private loss on client $i$’s local data
$\ell_{\text{pseudo}}=L(h_{i}^{t-1},P^{t})$ Loss on pseudo-labeled public data $P^{t}$
$L_{i}^{t}\in\{0,1\}^{|U|\times C}$ One-hot prediction matrix from client $i$ on public data
$E_{i}^{t}\in(0,\infty)^{|U|}$ Confidence (expertise) vector from client $i$ on public data
$S^{t}=\sum_{i=1}^{m}\text{diag}(E_{i}^{t})\cdot L_{i}^{t}$ Weighted score matrix used for consensus aggregation
$L^{t}[j]=\arg\max_{c\in[C]}S^{t}[j,c]$ Consensus pseudo-label for public example $x_{j}\in U$
Adaptive Weighting Mechanism
$\lambda_{i}^{t}$ Adaptive weight controlling trust in global signal for client $i$ at round $t$
$\ell=\ell_{\text{priv}}+\lambda_{i}^{t}\cdot\ell_{\text{pseudo}}$ Total loss used for local model update at round $t$
Optimization and Convergence
$\theta$ Model parameters
$\nabla L(\theta)$ Gradient of loss with respect to model parameters
$\sigma^{2}$ Bounded variance of local gradient estimator
$\tilde{\sigma}^{2}$ Bounded variance of global gradient estimator (pseudo-label noise)
$\delta$ Bounded drift in local objectives across rounds
$L$ Smoothness constant (Lipschitz constant of the gradient)
Sets and Indexing
$[m]=\{1,\dots,m\}$ Index set of all clients
$[C]=\{1,\dots,C\}$ Index set of all classes
$x_{j}\in U$ $j$-th public unlabeled sample
$y_{j}$ True (unknown) label of public sample $x_{j}$
$|U|$ Number of samples in the public dataset $U$
$|D_{i}|$ Number of samples in the local dataset of client $i$
$L_{i}^{t}[j,c]$ $(j,c)$-th entry of prediction matrix $L_{i}^{t}$
$E_{i}^{t}[j]$ Confidence of client $i$ on public example $x_{j}$